# Exploring Nonreciprocal Noise Transfer under Onsager-Casimir Symmetry in Synthetic-Field Optomechanics

### v1

@@ Abstract
An optomechanical system of fundamental importance consists of two intercoupled mechanical resonators, which are radiation-pressure coupled individually to a photonic cavity. This closed-loop and overall lossy configuration possesses two exceptional points (EPs) and offers the realization of synthetic magnetism, controlled by the loop phase. To elucidate the intricate role of loop phase and EPs in this setting, we analyze the noise power spectral density profiles of internal as well as output fluctuations. In the presence of a synthetic magnetic field, the non-reciprocal routing of a signal is well-known. Here, we further show that this also applies to non-reciprocal backaction noise flow when the time-reversal symmetry is broken, while the Onsager-Casimir symmetry still holds. To better quantify this phenomenon, we introduce a non-reciprocity measure that contrasts the time-reversed counterparts as a function of loop-phase. We observe that non-reciprocal noise flow is enhanced for smaller intermechanical couplings at the expense of lower sensitivity, whereas for sensing purposes, using a higher intermechanical coupling constant is the more viable option.
preprint: AIP/123-QED
@@ I Introduction
Open quantum systems described by non-Hermitian Hamiltonians with complex eigenvalue spectrum attract attention due to their aptness for modeling realistic systems in optics. El-Ganainy et al. (2018, 2019); Wang et al. (2023); Miri and Alu (2019); Özdemir et al. (2019) When a non-Hermitian system is steered in the parameter space, it can brought to an exceptional point (EP)- a non-Hermitian singularity with coalescing eigenvalues and eigenstates- that are special type of spectral singularities which open a venue for applications and technologies such as enhanced responsivity in optical and electromechanical sensors, Hodaei et al. (2017); Chen et al. (2017); Lu et al. (2021) realizing high-sensitivity measurements, Wu et al. (2023) optomechanically induced transparency, Sütlüoğlu and Bulutay (2021); Jing et al. (2015); Jiao et al. (2016) phonon lasing,Jing et al. (2014) adiabatic transport,Hoeller, Read, and Harris (2020) photonic quantum interference,Klauck et al. (2025) and enhanced cooling and squeezing. Sütlüoğlu Ege and Bulutay (2024) A particularly interesting example of non-Hermitian systems is unidirectional transport. Ramezani et al. (2010); Peng et al. (2014); Aleahmad et al. (2017) Such non-reciprocal and chiral effects have garnered significant attention in optics and photonics, Xu et al. (2016); Reisenbauer et al. (2024); Gou et al. (2020a); Choi et al. (2017); Laha et al. (2020); Lau and Clerk (2018); Lee et al. (2023); Soleymani et al. (2022) atomic and trapped ion systems Cai, Liu, and Yang (2024); Gou et al. (2020b); Bu et al. (2024); Zhang et al. (2025) and metamaterials. Buddhiraju et al. (2020); Lannebère et al. (2022); Schomerus (2020); Wang et al. (2021)
@@
A complementary, but conceptually rich framework for achieving non-reciprocity lies in the use of artificial gauge fields, especially in optomechanical systems where photons interact with mechanical vibrations, so that these systems become a suitable platform for simulating synthetic gauge fields. Schmidt et al. (2015); Walter and Marquardt (2016); Chen et al. (2021); Mathew, Pino, and Verhagen (2020); Ehrhardt et al. (2023) Interaction between mechanical and optical modes via radiation pressure allows phase-sensitive coupling, generating complex hopping amplitudes by tuning the phases of the external driving lasers. Chen et al. (2021) These gauge fields in the form of a complex phase factor introduce U(1) gauge symmetry in photonic and optomechanical systems.Aidelsburger, Nascimbene, and Goldman (2018) Introducing a synthetic gauge field with a nonzero phase around a plaquette breaks the time-reversal symmetry, leading to non-reciprocity in photonic and optomechanical systems. Hafezi and Rabl (2012); Manipatruni, Robinson, and Lipson (2009); Xu et al. (2020) Non-reciprocity enables robust, direction-dependent signal transport, which is essential for integrated photonic circuits, particularly in the implementation of nonreciprocal devices such as isolators, modulators, and circulators. Miri et al. (2017); Xu and Li (2015); Metelmann and Clerk (2015); Shen et al. (2016); Lee et al. (2025) Breaking reciprocity symmetries, inevitably opens additional noise pathways which couple into the measurement process.Xu, Liu, and Yu (2025); Ye et al. (2025)
@@
A significant obstacle in high-precision optical measurements is handling the introduced noise in the system. Optical systems’ performance of sensitivity contend two intertwined fundamental sources of uncertainty: one of them is induced with the fluctuations in the probe laser (imprecision or the so-called shot noise), and the other is the backaction noise introduced on the light while probing a system (e.g., radiation pressure noise in optomechanical systems). A primary challenge in quantum sensing is to identify the methods to circumvent both shot-noise and measurement backaction noise. Various techniques have been proposed to overcome those limitations, including the use of squeezed light, Lee, Lee, and Seok (2020); Purdy et al. (2013) backaction evading measurements schemes Shomroni et al. (2019); Hertzberg et al. (2010) and the coherent quantum noise cancellation. Tsang and Caves (2010); Wimmer et al. (2014); Bariani et al. (2015) While all these methods ultimately aim to enhance measurement sensitivity by alleviating noise, they operate through distinct physical mechanisms. A theoretical framework for ponderomotive squeezing of light through the radiation-pressure interaction between a laser field and a cryogenically cooled membrane achieved a 32% sensitivity below the shot-noise level. Purdy et al. (2013) Another method is directly injecting a coherent superposition of coherent light and squeezed vacuum field Lee, Lee, and Seok (2020) or frequency-dependent squeezed vacuum into a quantum sensing system. Subhash et al. (2022) While these squeezed light utilization methods reduce the imprecision noise by adjusting the input noise, alternative methods focus on backaction evading measurements by directly suppressing the radiation pressure noise. These methods utilize the fact that radiation pressure predominantly couples to a certain mechanical quadrature, allowing backaction evasion techniques to enable continuous measurement without disturbing the conjugate variable, and up to 0.67 dB (14%) reduction of total measurement noise can be achieved. Shomroni et al. (2019) Additionally, backaction noise across all frequencies can be eliminated via quantum coherent-noise cancellation by identifying interference pathways between quantum noises, Tsang and Caves (2010), adding an auxiliary optical cavity mimicking a negative-mass oscillator, Wimmer et al. (2014), or using an ensemble of ultracold atoms which provides a negative-mass dynamics. Bariani et al. (2015) Beyond these efforts for minimizing measurement noise, recent research focused on asymmetry in signal flow, particularly non-reciprocity, which is necessary for amplifiers and isolators. Interestingly, thermal-noise cancellation can be used for achieving optomechanically induced non-reciprocity when a whispering-gallery-mode cavity is coupled to mechanical resonators. The system can be designed so that thermal noise pathways from each mechanical resonator interfere destructively and cancel thermal noise in the system, culminating in asymmetric signal transmission and low thermal noise without needing pre-cooling mechanisms.Tang and Xu (2023)
@@
In this manuscript, we study theoretically a closed-loop optomechanical system consisting of a photonic cavity and intercoupled identical lossy mechanical resonators. This configuration, constituting a synthetic magnetism, enables a controllable non-reciprocal noise transfer mediated by the closed-loop phase. This is achieved when time-reversal symmetry is broken while the Onsager-Casimir relations are still valid.Brechet (2022) First, by exploring the intracavity and cavity output PSD, we demonstrate noise cancellation at the upper and lower bands depending on the loop phase. Next, we show that the internal noise PSD of mechanical resonators can be independently controlled, which paves the way for non-reciprocity. The most conspicuous outcome is the non-reciprocal backaction noise transfer between the mechanical resonators, and this flow is enhanced for smaller mechanical coupling constants. This optomechanical scheme can also be realized as a sensor with increased sensitivity beyond the first EP.
@@ II Theory
We consider a closed-loop optomechanical system consisting of a photonic cavity coupled to two lossy mechanical resonators via radiation pressure coupling constants $g_{1}$ and $g_{2}$. Loss rates of first and second mechanical resonators are $\gamma_{1}>0$ and $\gamma_{2}>0$, respectively. Both mechanical resonators share an equal resonance angular frequency $\omega_{m}$, and are intercoupled via mechanical coupling constant $\mu$. The photonic cavity is driven with a coherent pump laser with carrier frequency $\omega_{L}$ and amplitude $\varepsilon_{L}=\sqrt{\frac{2P_{L}}{\hbar\omega_{L}}}$, which is associated with the corresponding laser power $P_{L}$.
@@ II.1 Onsager-Casimir Relation
Onsager relations are established in the context of statistical physics, and in general, they assert that the linear response of a system is symmetric with respect to input and output. Specifically, when time-reversal symmetry and microscopic reversibility holds, Onsager symmetry guarantees that the linear response matrix entries, $L_{ij}$ have symmetric cross-couplings between paired fluxes and forces, i.e., $L_{ij}=L_{ji}$.Onsager (1931a, b) Onsager reciprocity relations are widely applicable in classical thermodynamics, and in electromagnetic systems are known under the name Lorentz reciprocity.Caloz et al. (2018) However, when time-reversal symmetry is broken, say under magnetic fluxes, synthetic fields, nonlinearity, or gain/loss engineering, these symmetry relations are generalized into Onsager-Casimir relations. They account for time-reversal parity of the associated quantities ($\varepsilon_{i}=+1$ for even, and $\varepsilon_{i}=-1$ for odd parity), and impose a modified symmetry $L_{ij}(B)=\varepsilon_{i}\varepsilon_{j}L_{ji}(-B)$ which mandates the system’s dynamics remain microscopically reversible under magnetic field $(B)$ reversal.Brechet (2022) This generalization forms the theoretical foundation for understanding and designing non-Hermitian and non-reciprocal photonic devices. Tsang (2024); Vicencio et al. (2024) For our optomechanical system, where the closed-loop coupling phase with U(1) symmetry introduces synthetic magnetism, the generalized Onsager-Casimir relations hold. In other words, the response is invariant if the mechanical resonators are interchanged ($1\leftrightarrow 2$) together with negating the loop phase, i.e., $\phi_{\ell}\rightarrow-\phi_{\ell}$.
@@ II.3 Quantum Langevin equations
$+\eta\kappa\varepsilon_{L}-\frac{\kappa}{2}\hat{a}+\sqrt{\kappa}\hat{a}_{in}(t),$
$\frac{d\hat{b}_{1}}{dt}$ $=$ $-i\omega_{m}\hat{b}_{1}+i\mu\hat{b}_{2}+ig_{1}\hat{a}^{\dagger}\hat{a}-\frac{\gamma_{1}}{2}\hat{b}_{1}+\sqrt{\gamma}_{1}\hat{b}_{1,in}(t),$
$\frac{d\hat{b}_{2}}{dt}$ $=$ $-i\omega_{m}\hat{b}_{2}+i\mu^{*}\hat{b}_{1}+ig_{2}\hat{a}^{\dagger}\hat{a}-\frac{\gamma_{2}}{2}\hat{b}_{2}+\sqrt{\gamma}_{2}\hat{b}_{2,in}(t),$
where $\hat{a}_{in}(t)$, $\hat{b}_{1,in}(t)$, and $\hat{b}_{2,in}(t)$ are zero-mean cavity input noise and mechanical thermal noise operators, respectively.Gardiner and Zoller (2004) Photonic and mechanical Markovian input noise operators satisfy the following non-zero correlations
$\langle\hat{a}_{in}^{\dagger}(t)\hat{a}_{in}(t^{\prime})\rangle$ $=$ $n_{a}\delta(t-t^{\prime}),$ (2)
$\langle\hat{a}_{in}(t)\hat{a}_{in}^{\dagger}(t^{\prime})\rangle$ $=$ $(n_{a}+1)\delta(t-t^{\prime}),$
$\langle\hat{b}_{j,in}^{\dagger}(t)\hat{b}_{j,in}(t^{\prime})\rangle$ $=$ $n_{b_{j}}\delta(t-t^{\prime}),$
$\langle\hat{b}_{j,in}(t)\hat{b}_{j,in}^{\dagger}(t^{\prime})\rangle$ $=$ $(n_{b_{j}}+1)\delta(t-t^{\prime}),$
for $j=1,2$, and where $n_{a}$ and $n_{b_{j}}$ are the mean thermal occupancies of the optical cavity and respective thermal reservoirs.Clerk et al. (2010) Since our main focus is on fluctuations around the steady-state mean values, subsequently we linearize quantum Langevin equations by expanding each mode as a sum of the classical mean value and quantum fluctuation, $\hat{\aleph}(t)\rightarrow\langle\hat{\aleph}(t)\rangle+\delta\hat{\aleph}(t)$ where $\hat{\aleph}=\hat{a},\hat{b}_{1},\hat{b}_{2}$. The steady-state mean values are given by
@@
$\frac{d\delta\hat{a}}{dt}$ $=$ $i(-\Delta_{a}+i\kappa/2)\delta\hat{a}+\sqrt{\kappa}\hat{a}_{in}(t)+i\langle\hat{a}\rangle g_{1}^{*}\delta\hat{b}_{1}$
$+i\langle\hat{a}\rangle g_{1}\delta\hat{b}_{1}^{\dagger}+i\langle\hat{a}\rangle g_{2}^{*}\delta\hat{b}_{2}+i\langle\hat{a}\rangle g_{2}\delta\hat{b}_{2}^{\dagger},$
@@
$+ig_{1}\langle\hat{a}\rangle\delta\hat{a}^{\dagger}+\sqrt{\gamma_{1}}\hat{b}_{1,in}(t),$
@@
$+ig_{2}\langle\hat{a}\rangle\delta\hat{a}^{\dagger}+\sqrt{\gamma_{2}}\hat{b}_{2,in}(t).$
@@
$\hat{X}^{in}_{\aleph=a,b_{1},b_{2}}$ $=$ $\frac{\hat{\aleph}_{in}+\hat{\aleph}_{in}^{\dagger}}{\sqrt{2}},$
$\hat{Y}^{in}_{\aleph=a,b_{1},b_{2}}$ $=$ $\frac{\hat{\aleph}_{in}-\hat{\aleph}_{in}^{\dagger}}{i\sqrt{2}}.$
@@
where over dot represents time derivative, and $\mathbf{\hat{R}}(t)=[\delta\hat{X}_{a},\delta\hat{Y}_{a},\delta\hat{X}_{b_{1}},\delta\hat{Y}_{b_{1}},\delta\hat{X}_{b_{2}},\delta\hat{Y}_{b_{2}}]^{T}$ stands for quadrature fluctuation operators, and $\mathbf{M}$ is a $\times6$ time-independent matrix which consists of quantum fluctuation coefficients. Noise vector $\mathbf{\hat{N}}(t)$ is defined as the product of the input noise vector, $\mathbf{\hat{R}}^{in}(t)$, and a diagonal coefficient matrix $\mathbf{D}$, which are given as,
$\mathbf{\hat{R}}^{in}(t)=\left[\hat{X}^{in}_{a},\hat{Y}^{in}_{a},\hat{X}^{in}_{b_{1}},\hat{Y}^{in}_{b_{1}},\hat{X}^{in}_{b_{2}},\hat{Y}^{in}_{b_{2}}\right]^{T},$
and
$\mathbf{D}=\begin{pmatrix}\sqrt{\kappa}&0&0&0&0&0\\ 0&\sqrt{\kappa}&0&0&0&0\\ 0&0&\sqrt{\gamma_{1}}&0&0&0\\ 0&0&0&\sqrt{\gamma_{1}}&0&0\\ 0&0&0&0&\sqrt{\gamma_{2}}&0\\ 0&0&0&0&0&\sqrt{\gamma_{2}}\\ \end{pmatrix}$.

### v2

@@ Abstract
An optomechanical system of fundamental importance consists of two intercoupled mechanical resonators, which are radiation-pressure coupled individually to a photonic cavity. This closed-loop and overall lossy configuration possesses two exceptional points (EPs) and offers the realization of synthetic magnetism, controlled by the loop phase. To elucidate the intricate role of loop phase and EPs in this setting, we analyze the noise power spectral density profiles of internal as well as output fluctuations. In the presence of a synthetic magnetic field, the nonreciprocal routing of a signal is well known. Here, we further show that this also applies to nonreciprocal backaction noise flow when the time-reversal symmetry is broken, while the Onsager-Casimir symmetry still holds. To better quantify this phenomenon, we introduce a nonreciprocity measure that contrasts the time-reversed counterparts as a function of loop phase. We observe that nonreciprocal noise flow is enhanced for smaller intermechanical couplings at the expense of lower sensitivity, whereas for sensing purposes, using a higher intermechanical coupling constant is the more viable option.
@@ I Introduction
Open quantum systems described by non-Hermitian Hamiltonians with complex eigenvalue spectrum attract attention due to their aptness for modeling realistic systems in optics [1, 2, 3, 4, 5]. When a non-Hermitian system is steered in the parameter space, it can brought to an exceptional point (EP)—a non-Hermitian singularity with coalescing eigenvalues and eigenstates—that are special type of spectral singularities which open a venue for applications and technologies such as enhanced responsivity in optical and electromechanical sensors [6, 7, 8], realizing high-sensitivity measurements [9], optomechanically induced transparency [10, 11, 12], phonon lasing [13], adiabatic transport [14], photonic quantum interference [15], and enhanced cooling and squeezing [16]. A particularly interesting example of non-Hermitian systems is unidirectional transport [17, 18, 19]. Such nonreciprocal and chiral effects have garnered significant attention in optics and photonics [20, 21, 22, 23, 24, 25, 26, 27], atomic and trapped-ion systems [28, 29, 30], and metamaterials [31, 32, 33, 34].
@@
A complementary, but conceptually rich framework for achieving nonreciprocity invokes the use of artificial gauge fields, especially in optomechanical systems where photons interact with mechanical vibrations, so that these systems become a suitable platform for simulating synthetic gauge fields [35, 36, 37, 38, 39]. Interaction between mechanical and optical modes via radiation pressure allows phase-sensitive coupling, generating complex hopping amplitudes by tuning the phases of the external driving lasers [37]. These gauge fields in the form of a complex phase factor introduce U(1) gauge symmetry in photonic and optomechanical systems [40]. Introducing a synthetic gauge field with a nonzero phase around a plaquette breaks the time-reversal symmetry, leading to nonreciprocity in photonic and optomechanical systems [41, 42, 43]. Nonreciprocity enables robust, direction-dependent signal transport, which is essential for integrated photonic circuits, particularly in the implementation of nonreciprocal devices such as isolators, modulators, and circulators [44, 45, 46, 47, 48]. Breaking reciprocity symmetries, inevitably opens additional noise pathways which couple into the measurement process [49, 50]. Building on the discussions of $U(1)$ gauge symmetry and artificial gauge fields that are the key mechanisms for nonreciprocity, a fundamental concept is the Onsager-Casimir relation. This is the generalization of the Onsager relation, which states that linear response of a system near equilibrium is symmetric under time-reversal symmetry and microscopic reversibility [51, 52]. This relation turns into Onsager-Casimir relation when time-reversal symmetry is broken [53].
@@
A significant obstacle in high-precision optical measurements is handling the introduced noise in the system. Optical systems’ performance of sensitivity contend two intertwined fundamental sources of uncertainty: One of them is induced with the fluctuations in the probe laser (imprecision or the so-called shot noise), and the other is the backaction noise introduced on the light while probing a system (e.g., radiation-pressure noise in optomechanical systems). A primary challenge in quantum sensing is to identify the methods to circumvent both shot noise and measurement backaction noise. Various techniques have been proposed to overcome those limitations, including the use of squeezed light, [54, 55] backaction-evading measurements schemes [56, 57], and the coherent quantum noise cancellation [58, 59, 60]. While all these methods ultimately aim to enhance measurement sensitivity by alleviating noise, they operate through distinct physical mechanisms. A theoretical framework for ponderomotive squeezing of light through the radiation-pressure interaction between a laser field and a cryogenically cooled membrane achieved a 32% sensitivity below the shot-noise level [55]. Another method is directly injecting a coherent superposition of coherent light and squeezed vacuum field [54] or frequency-dependent squeezed vacuum into a quantum sensing system [61]. While these squeezed-light utilization methods reduce the imprecision noise by adjusting the input noise, alternative methods focus on backaction-evading measurements by directly suppressing the radiation-pressure noise. These methods utilize the fact that radiation pressure predominantly couples to a certain mechanical quadrature, allowing backaction evasion techniques to enable continuous measurement without disturbing the conjugate variable, and up to 0.67 dB (14%) reduction of total measurement noise can be achieved [56]. Additionally, backaction noise across all frequencies can be eliminated via quantum coherent-noise cancellation by identifying interference pathways between quantum noises [58], adding an auxiliary optical cavity mimicking a negative-mass oscillator [59], or using an ensemble of ultracold atoms, which provides a negative-mass dynamics [60]. Beyond these efforts for minimizing measurement noise, recent research focused on asymmetry in signal flow, particularly nonreciprocity, which is necessary for amplifiers and isolators. Interestingly, thermal-noise cancellation can be used for achieving optomechanically induced nonreciprocity when a whispering-gallery-mode cavity is coupled to mechanical resonators. The system can be designed so that thermal-noise pathways from each mechanical resonator interfere destructively and cancel thermal noise in the system, culminating in asymmetric signal transmission and low thermal noise without needing precooling mechanisms [62].
@@
In this article, we study theoretically a closed-loop optomechanical system consisting of a photonic cavity and intercoupled identical lossy mechanical resonators. This configuration, constituting a synthetic magnetism, enables a controllable nonreciprocal noise transfer mediated by the closed-loop phase. This is achieved when time-reversal symmetry is broken while the Onsager-Casimir relation is still valid [53]. First, by exploring the intracavity and cavity-output power spectral density (PSD), we demonstrate noise cancellation at the upper and lower bands depending on the loop phase. Next, we show that the internal noise PSD of mechanical resonators can be independently controlled, which paves the way for nonreciprocity. The most conspicuous outcome is the nonreciprocal backaction noise transfer between the mechanical resonators, and this flow is enhanced for smaller mechanical coupling constants. This optomechanical scheme can also be realized as a sensor with increased sensitivity beyond the first EP.
@@ II Theory
We consider a closed-loop optomechanical system consisting of a photonic cavity coupled to two lossy mechanical resonators via radiation-pressure coupling constants $g_{1}$ and $g_{2}$ as shown in Fig. 1. Loss rates of first and second mechanical resonators are $\gamma_{1}>0$ and $\gamma_{2}>0$, respectively. Both mechanical resonators share an equal resonance angular frequency $\omega_{m}$, and are intercoupled via mechanical coupling constant $\mu$. The photonic cavity is driven with a coherent pump laser with carrier frequency $\omega_{L}$ and amplitude $\varepsilon_{L}=\sqrt{\frac{2P_{L}}{\hbar\omega_{L}}}$, which is associated with the corresponding laser power $P_{L}$.
@@ II.1 Onsager-Casimir relation
Onsager relations are established in the context of statistical physics, and in general, they assert that the linear response of a system is symmetric with respect to input and output. Specifically, when time-reversal symmetry and microscopic reversibility holds, Onsager symmetry guarantees that the linear response matrix entries, $L_{ij}$ have symmetric cross-couplings between paired fluxes and forces, i.e., $L_{ij}=L_{ji}$ [51, 52]. Onsager reciprocity relations are widely applicable in classical thermodynamics, and in electromagnetic systems are known under the name Lorentz reciprocity [63]. However, when time-reversal symmetry is broken, say under magnetic fluxes, synthetic fields, nonlinearity, or gain/loss engineering, these symmetry relations are generalized into Onsager-Casimir relations. They account for time-reversal parity of the associated quantities ($\varepsilon_{i}=+1$ for even, and $\varepsilon_{i}=-1$ for odd parity), and impose a modified symmetry $L_{ij}(B)=\varepsilon_{i}\varepsilon_{j}L_{ji}(-B)$, which mandates that the system’s dynamics remain microscopically reversible under magnetic field $(B)$ reversal [53]. This generalization forms the theoretical foundation for understanding and designing non-Hermitian and nonreciprocal photonic devices [64, 65]. For our optomechanical system, where the closed-loop coupling phase with $U(1)$ symmetry introduces synthetic magnetism, the generalized Onsager-Casimir relations hold. In other words, the response is invariant if the mechanical resonators are interchanged ($1\leftrightarrow 2$) together with negating the loop phase, i.e., $\phi_{\ell}\rightarrow-\phi_{\ell}$.
@@ II.3 Quantum Langevin equations
$+\eta\kappa\varepsilon_{L}-\frac{\kappa}{2}\hat{a}+\sqrt{\kappa}\hat{a}_{\mathrm{in}}(t),$
$\frac{d\hat{b}_{1}}{dt}$ $=$ $-i\omega_{m}\hat{b}_{1}+i\mu\hat{b}_{2}+ig_{1}\hat{a}^{\dagger}\hat{a}-\frac{\gamma_{1}}{2}\hat{b}_{1}+\sqrt{\gamma}_{1}\hat{b}_{1,\mathrm{in}}(t),$
$\frac{d\hat{b}_{2}}{dt}$ $=$ $-i\omega_{m}\hat{b}_{2}+i\mu^{*}\hat{b}_{1}+ig_{2}\hat{a}^{\dagger}\hat{a}-\frac{\gamma_{2}}{2}\hat{b}_{2}+\sqrt{\gamma}_{2}\hat{b}_{2,\mathrm{in}}(t),$
where $\hat{a}_{\mathrm{in}}(t)$, $\hat{b}_{1,\mathrm{in}}(t)$, and $\hat{b}_{2,\mathrm{in}}(t)$ are zero-mean cavity input noise and mechanical thermal noise operators, respectively [66]. Photonic and mechanical Markovian input noise operators satisfy the following non-zero correlations
$\langle\hat{a}_{\mathrm{in}}^{\dagger}(t)\hat{a}_{\mathrm{in}}(t^{\prime})\rangle$ $=$ $n_{a}\delta(t-t^{\prime}),$ (1)
$\langle\hat{a}_{\mathrm{in}}(t)\hat{a}_{\mathrm{in}}^{\dagger}(t^{\prime})\rangle$ $=$ $(n_{a}+1)\delta(t-t^{\prime}),$
$\langle\hat{b}_{j,\mathrm{in}}^{\dagger}(t)\hat{b}_{j,\mathrm{in}}(t^{\prime})\rangle$ $=$ $n_{b_{j}}\delta(t-t^{\prime}),$
$\langle\hat{b}_{j,\mathrm{in}}(t)\hat{b}_{j,\mathrm{in}}^{\dagger}(t^{\prime})\rangle$ $=$ $(n_{b_{j}}+1)\delta(t-t^{\prime}),$
for $j=1,2$, and where $n_{a}$ and $n_{b_{j}}$ are the mean thermal occupancies of the optical cavity and respective thermal reservoirs [67]. Since our main focus is on fluctuations around the steady-state mean values, subsequently we linearize quantum Langevin equations by expanding each mode as a sum of the classical mean value and quantum fluctuation, $\hat{\aleph}(t)\rightarrow\langle\hat{\aleph}(t)\rangle+\delta\hat{\aleph}(t)$, where $\hat{\aleph}=\hat{a},\hat{b}_{1},\hat{b}_{2}$. The steady-state mean values are given by
@@
$\frac{d\delta\hat{a}}{dt}$ $=$ $i(-\Delta_{a}+i\kappa/2)\delta\hat{a}+\sqrt{\kappa}\hat{a}_{\mathrm{in}}(t)$
$+i\langle\hat{a}\rangle g_{1}^{*}\delta\hat{b}_{1}+i\langle\hat{a}\rangle g_{1}\delta\hat{b}_{1}^{\dagger}+i\langle\hat{a}\rangle g_{2}^{*}\delta\hat{b}_{2}+i\langle\hat{a}\rangle g_{2}\delta\hat{b}_{2}^{\dagger},$
@@
$+ig_{1}\langle\hat{a}\rangle\delta\hat{a}^{\dagger}+\sqrt{\gamma_{1}}\hat{b}_{1,\mathrm{in}}(t),$
@@
$+ig_{2}\langle\hat{a}\rangle\delta\hat{a}^{\dagger}+\sqrt{\gamma_{2}}\hat{b}_{2,\mathrm{in}}(t).$
@@
$\hat{X}^{\mathrm{in}}_{\aleph=a,b_{1},b_{2}}$ $=$ $\frac{\hat{\aleph}_{\mathrm{in}}+\hat{\aleph}_{\mathrm{in}}^{\dagger}}{\sqrt{2}},$
$\hat{Y}^{\mathrm{in}}_{\aleph=a,b_{1},b_{2}}$ $=$ $\frac{\hat{\aleph}_{\mathrm{in}}-\hat{\aleph}_{\mathrm{in}}^{\dagger}}{i\sqrt{2}}.$
@@
where over dot represents time derivative, and $\mathbf{\hat{R}}(t)=[\delta\hat{X}_{a},\delta\hat{Y}_{a},\delta\hat{X}_{b_{1}},\delta\hat{Y}_{b_{1}},\delta\hat{X}_{b_{2}},\delta\hat{Y}_{b_{2}}]^{T}$ stands for quadrature fluctuation operators, and $\mathbf{M}$ is a $\times6$ time-independent drift matrix, which consists of quantum fluctuation coefficients
[…]
where for notational convenience we define $\Xi\coloneq\Delta-2(\operatorname{Re}(g_{1})\operatorname{Re}(\langle\hat{b}_{1}\rangle)+\operatorname{Re}(g_{2})\operatorname{Re}(\langle\hat{b}_{2}\rangle))-2(\operatorname{Im}(g_{1})\operatorname{Im}(\langle\hat{b}_{1}\rangle)+\operatorname{Im}(g_{2})\operatorname{Im}(\langle\hat{b}_{2}\rangle))$. Noise vector $\mathbf{\hat{N}}(t)$ is defined as
$\mathbf{\hat{N}}(t)=\big[$ $\sqrt{\kappa}\hat{X}^{\mathrm{in}}_{a},\sqrt{\kappa}\hat{Y}^{\mathrm{in}}_{a},\sqrt{\gamma_{1}}\hat{X}^{\mathrm{in}}_{b_{1}},\sqrt{\gamma_{1}}\hat{Y}^{\mathrm{in}}_{b_{1}},\sqrt{\gamma_{2}}\hat{X}^{\mathrm{in}}_{b_{2}},$ (4)
$\sqrt{\gamma_{2}}\hat{Y}^{\mathrm{in}}_{b_{2}}\big]^{T},$
which is the product of the input noise vector, $\mathbf{\hat{R}}^{\mathrm{in}}(t)$,
$\mathbf{\hat{R}}^{in}(t)=\left[\hat{X}^{\mathrm{in}}_{a},\hat{Y}^{\mathrm{in}}_{a},\hat{X}^{\mathrm{in}}_{b_{1}},\hat{Y}^{\mathrm{in}}_{b_{1}},\hat{X}^{in}_{b_{2}},\hat{Y}^{\mathrm{\mathrm{in}}}_{b_{2}}\right]^{T},$ (5)
and a diagonal coefficient matrix $\mathbf{D}$ as
$\mathbf{D}=\mathrm{diag}\left[\sqrt{\kappa},\sqrt{\kappa},\sqrt{\gamma_{1}},\sqrt{\gamma_{1}},\sqrt{\gamma_{2}},\sqrt{\gamma_{2}}\right],$ (6)
whose diagonal entries represent the individual loss rates associated with each resonator.